# Latent Schrödinger Bridge Diffusion Model for Generative Learning

### v1

@@ C.2 Statistical Error
[…]
$\leq$ $\mathop{\mathrm{sup}}_{\mathbf{s}\in\mathrm{NN}}\left(\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{z}}h_{\mathbf{s}}(t,\widehat{\mathbf{x}}_{i},\mathbf{z})\mathrm{d}t-\frac{1}{m}\sum_{j=1}^{m}h_{\mathbf{s}}(t_{j},\widehat{\mathbf{x}}_{i},\mathbf{z}_{j})\right)$
$\leq$ $\mathop{\mathrm{max}}_{\mathbf{s}_{\delta}\in\mathcal{C}_{\delta}}\left(\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{z}}h_{\mathbf{s}_{\delta}}(t,\mathbf{x}_{i},\mathbf{z})\mathrm{d}t-\frac{1}{m}\sum_{j=1}^{m}h_{\mathbf{s}_{\delta}}(t_{j},\mathbf{x}_{i},\mathbf{z}_{j})\right)+4\left(K+\frac{\sqrt{d^{*}}r_{0}}{\sigma\sqrt{1-T}}\right)\delta.$
Let $b=4\left(K+\frac{\sqrt{d^{*}}r_{0}}{\sigma\sqrt{1-T}}\right)\delta$. For $t>b$, using Hoeffding’s inequality implies
@@
$\leq$ $\leavevmode\nobreak\ \mathbb{P}_{\mathcal{T},\mathcal{Z}}\left(\mathop{\mathrm{max}}_{\mathbf{s}_{\delta}\in\mathcal{C}_{\delta}}\left(\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{z}}h_{\mathbf{s}_{\delta}}(t,\widehat{\mathbf{x}}_{i},\mathbf{z})\mathrm{d}t-\frac{1}{m}\sum_{j=1}^{m}h_{\mathbf{s}_{\delta}}(t_{j},\widehat{\mathbf{x}}_{i},\mathbf{z}_{j})\right)>t-b\right)$
$\leq$ $\leavevmode\nobreak\ \mathcal{N}\mathop{\mathrm{max}}_{\mathbf{s}_{\delta}\in\mathcal{C}_{\delta}}\mathbb{P}_{\mathcal{T},\mathcal{Z}}\left(\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{z}}h_{\mathbf{s}_{\delta}}(t,\widehat{\mathbf{x}}_{i},\mathbf{z})\mathrm{d}t-\frac{1}{m}\sum_{j=1}^{m}h_{\mathbf{s}_{\delta}}(t_{j},\widehat{\mathbf{x}}_{i},\mathbf{z}_{j})>t-b\right)$
$\leq$ $\leavevmode\nobreak\ \mathcal{N}\exp{\left(-\frac{2m(t-b)^{2}}{E_{T}^{2}(r_{0})}\right)}.$
Therefore, by taking expectation over $\mathcal{T},\mathcal{Z}$, we deduce that $(B)$ satisfies
@@
$\leq b+c_{0}+\mathcal{N}\int_{c_{0}}^{+\infty}\exp{\left(-\frac{2mt^{2}}{E_{T}^{2}(r_{0})}\right)}\mathrm{d}t$
$\leq b+c_{0}+\frac{\sqrt{\pi}}{2}\mathcal{N}\exp{\left(-\frac{2mc_{0}^{2}}{E_{T}^{2}(r_{0})}\right)}\frac{E_{T}(r_{0})}{\sqrt{2m}}.$
@@
$\leq$ $\leavevmode\nobreak\ b+c_{0}+\frac{\sqrt{\pi}}{2}\mathcal{N}\exp{\left(-\frac{2mc_{0}^{2}}{E_{T}^{2}(r_{0})}\right)}\frac{E_{T}(r_{0})}{\sqrt{2m}}.$
@@
$(C)=-\frac{1}{mn}\sum_{i=1}^{n}\sum_{j=1}^{m}\left\|\widehat{\mathbf{s}}(t_{j},\widehat{\mathbf{x}}_{i}+\sigma\sqrt{1-t_{j}}\mathbf{z}_{j})+\frac{\mathbf{z}_{j}}{\sigma\sqrt{1-t_{j}}}\right\|^{2}\mathbf{I}_{\{\|\mathbf{z}_{j}\|_{\infty}>r_{0}\}}\leq 0,$
@@
$\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left[\frac{1}{n}\sum_{i=1}^{n}\left(\ell_{\widehat{\mathbf{s}}}(\widehat{\mathbf{x}}_{i})-\widehat{\ell}_{\widehat{\mathbf{s}}}(\widehat{\mathbf{x}}_{i})\right)\right]\leq b_{0}+c_{0}+\frac{\sqrt{\pi}}{2}\mathcal{N}\exp{\left(-\frac{2mc_{0}^{2}}{E_{T}^{2}(r_{0})}\right)}\frac{E_{T}(r_{0})}{\sqrt{2m}},$
where $b_{0}=\left(2K^{2}-\frac{2\log{(1-T)}\sqrt{d^{*}(d^{*}+2)}}{\sigma^{2}T}\right)\sqrt{2d^{*}}\exp\left(-\frac{r_{0}^{2}}{4}\right)+4\left(K+\frac{\sqrt{d^{*}}r_{0}}{\sigma\sqrt{1-T}}\right)\delta$.
Setting $c_{0}=E_{T}(r_{0})\sqrt{\frac{\log{\mathcal{N}}}{2m}}$, $r_{0}=2\sqrt{\log{m}}$, and $\delta=\frac{1}{m}$, we obtain
$\leavevmode\nobreak\ \mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left[\frac{1}{n}\sum_{i=1}^{n}\left(\ell_{\widehat{\mathbf{s}}}(\widehat{\mathbf{x}}_{i})-\widehat{\ell}_{\widehat{\mathbf{s}}}(\widehat{\mathbf{x}}_{i})\right)\right]$
$\leq$ $\leavevmode\nobreak\ b_{0}+E_{T}(r_{0})\cdot\frac{\sqrt{\log\mathcal{N}}+1}{\sqrt{2m}}$
$\leq$ $\leavevmode\nobreak\ \frac{\left(2K^{2}-\frac{2\log{(1-T)}\sqrt{d^{*}(d^{*}+2)}}{\sigma^{2}T}\right)\sqrt{2d^{*}}+4\left(K+\frac{2\sqrt{d^{*}}\sqrt{\log{m}}}{\sigma\sqrt{1-T}}\right)}{m}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(2K^{2}+\frac{4d^{*}\log{m}}{\sigma^{2}(1-T)}\right)\cdot\frac{\sqrt{\log\mathcal{N}}+1}{\sqrt{2m}}$
$=$ $\leavevmode\nobreak\ \widetilde{\mathcal{O}}\left(\frac{K^{2}}{m}+\frac{K^{2}}{\sqrt{m}}\cdot\frac{\xi^{\frac{d}{2}}\epsilon^{-\frac{d^{*}+1}{2}}}{(1-T)^{\frac{3}{2}}}\right)$
@@
C.3 Error Bound for Score Estimation

### v2

@@ Appendix C Statistical Error
$\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{z}}h_{\widehat{\mathbf{s}}}(t,\widehat{\mathbf{x}}_{i},\mathbf{z})\mathrm{d}t-\frac{1}{m}\sum_{j=1}^{m}h_{\widehat{\mathbf{s}}}(t_{j},\mathbf{x}_{i},\mathbf{z}_{j})$
$\leq$ $\mathop{\mathrm{sup}}_{\mathbf{s}\in\mathcal{C}}\left(\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{z}}h_{\mathbf{s}}(t,\widehat{\mathbf{x}}_{i},\mathbf{z})\mathrm{d}t-\frac{1}{m}\sum_{j=1}^{m}h_{\mathbf{s}}(t_{j},\widehat{\mathbf{x}}_{i},\mathbf{z}_{j})\right)$
$\leq$ $\mathop{\mathrm{max}}_{\mathbf{s}_{\delta}\in\mathcal{C}_{\delta_{1}}}\left(\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{z}}h_{\mathbf{s}_{\delta_{1}}}(t,\mathbf{x}_{i},\mathbf{z})\mathrm{d}t-\frac{1}{m}\sum_{j=1}^{m}h_{\mathbf{s}_{\delta_{1}}}(t_{j},\mathbf{x}_{i},\mathbf{z}_{j})\right)+2C_{12}{\mathcal{R}}^{\frac{C_{T}}{2}}(r+\sqrt{\log\mathcal{R}})\delta_{1}.$
Let $b=2C_{12}{\mathcal{R}}^{\frac{C_{T}}{2}}(r+\sqrt{\log\mathcal{R}})\delta_{1}$. For $t>b$, using Hoeffding’s inequality implies
@@
$\leq$ $\leavevmode\nobreak\ \mathbb{P}_{\mathcal{T},\mathcal{Z}}\left(\mathop{\mathrm{max}}_{\mathbf{s}_{\delta_{1}}\in\mathcal{C}_{\delta_{1}}}\left(\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{z}}h_{\mathbf{s}_{\delta_{1}}}(t,\widehat{\mathbf{x}}_{i},\mathbf{z})\mathrm{d}t-\frac{1}{m}\sum_{j=1}^{m}h_{\mathbf{s}_{\delta_{1}}}(t_{j},\widehat{\mathbf{x}}_{i},\mathbf{z}_{j})\right)>t-b\right)$
$\leq$ $\leavevmode\nobreak\ \mathcal{N}_{\delta_{1}}\mathop{\mathrm{max}}_{\mathbf{s}_{\delta_{1}}\in\mathcal{C}_{\delta_{1}}}\mathbb{P}_{\mathcal{T},\mathcal{Z}}\left(\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{z}}h_{\mathbf{s}_{\delta_{1}}}(t,\widehat{\mathbf{x}}_{i},\mathbf{z})\mathrm{d}t-\frac{1}{m}\sum_{j=1}^{m}h_{\mathbf{s}_{\delta_{1}}}(t_{j},\widehat{\mathbf{x}}_{i},\mathbf{z}_{j})>t-b\right)$
$\leq$ $\leavevmode\nobreak\ \mathcal{N}_{\delta_{1}}\exp{\left(-\frac{2m(t-b)^{2}}{E_{\mathcal{R}}^{2}(r)}\right)}.$
Therefore, by taking expectation over $\mathcal{T},\mathcal{Z}$, for any $c_{0}>0$, we deduce that $(B)$ satisfies
@@
$\leq b+c_{0}+\mathcal{N}_{\delta_{1}}\int_{c_{0}}^{+\infty}\exp{\left(-\frac{2mt^{2}}{E_{\mathcal{R}}^{2}(r)}\right)}\mathrm{d}t$
$\leq b+c_{0}+\frac{\sqrt{\pi}}{2}\mathcal{N}_{\delta_{1}}\exp{\left(-\frac{2mc_{0}^{2}}{E_{\mathcal{R}}^{2}(r)}\right)}\frac{E_{\mathcal{R}}(r)}{\sqrt{2m}}.$
@@
$\leq$ $\leavevmode\nobreak\ b+c_{0}+\frac{\sqrt{\pi}}{2}\mathcal{N}_{\delta_{1}}\exp{\left(-\frac{2mc_{0}^{2}}{E_{\mathcal{R}}^{2}(r)}\right)}\frac{E_{\mathcal{R}}(r)}{\sqrt{2m}}.$
@@
$(C)=-\frac{1}{mn}\sum_{i=1}^{n}\sum_{j=1}^{m}\left\|\widehat{\mathbf{s}}(1-t_{j},\widehat{\mathbf{x}}_{i}+\sigma\sqrt{1-t_{j}}\mathbf{z}_{j})+\frac{\mathbf{z}_{j}}{\sigma\sqrt{1-t_{j}}}\right\|^{2}\mathbf{I}_{\{\|\mathbf{z}_{j}\|_{\infty}>r\}}\leq 0,$
@@
$\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left[\frac{1}{n}\sum_{i=1}^{n}\left(\ell_{\widehat{\mathbf{s}}}(\widehat{\mathbf{x}}_{i})-\widehat{\ell}_{\widehat{\mathbf{s}}}(\widehat{\mathbf{x}}_{i})\right)\right]\leq b_{0}+c_{0}+\frac{\sqrt{\pi}}{2}\mathcal{N}_{\delta_{1}}\exp{\left(-\frac{2mc_{0}^{2}}{E_{\mathcal{R}}^{2}(r)}\right)}\frac{E_{\mathcal{R}}(r)}{\sqrt{2m}},$
where $b_{0}=C_{10}\log^{2}\mathcal{R}\exp\left(-\frac{r^{2}}{4}\right)+2C_{12}{\mathcal{R}}^{\frac{C_{T}}{2}}(r+\sqrt{\log\mathcal{R}})\delta_{1}$. By setting $c_{0}=E_{\mathcal{R}}(r)\sqrt{\frac{\log{\mathcal{N}_{\delta_{1}}}}{2m}}$, $r=2\sqrt{\log{m}}$, and $\delta_{1}=\frac{1}{m}$, we obtain
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left[\frac{1}{n}\sum_{i=1}^{n}\left(\ell_{\widehat{\mathbf{s}}}(\widehat{\mathbf{x}}_{i})-\widehat{\ell}_{\widehat{\mathbf{s}}}(\widehat{\mathbf{x}}_{i})\right)\right]$ (22)
$\leq b_{0}+E_{\mathcal{R}}(r)\cdot\frac{\sqrt{\log\mathcal{N}_{1/m}}+1}{\sqrt{2m}}$
$\lesssim\frac{\log^{2}\mathcal{R}+{\mathcal{R}}^{\frac{C_{T}}{2}}(\sqrt{\log m}+\sqrt{\log\mathcal{R}})}{m}+{\mathcal{R}}^{C_{T}}(\log m+\log\mathcal{R})\cdot\frac{{\mathcal{R}}^{\frac{d^{*}}{2}}\log^{\frac{13}{2}}\mathcal{R}(\log^{2}\mathcal{R}+\sqrt{\log m})}{\sqrt{m}}$
$\lesssim{\mathcal{R}}^{C_{T}}(\log m+\log\mathcal{R})\cdot\frac{{\mathcal{R}}^{\frac{d^{*}}{2}}\log^{\frac{13}{2}}\mathcal{R}(\log^{2}\mathcal{R}+\sqrt{\log m})}{\sqrt{m}}.$
The proof is complete. ∎
Combining Lemma 4.3 and Lemma 4.4, we can prove Theorem 4.5.
Proof of Theorem 4.5.
By choosing $\mathcal{R}=\lfloor n^{\frac{1}{d^{*}+2\beta}}\rfloor+1\lesssim n^{\frac{1}{d^{*}+2\beta}}$ in Lemma 4.3, the approximation error can be bounded as
$\inf_{\mathbf{s}\in\mathcal{C}}\left(\mathcal{L}(\mathbf{s})-\mathcal{L}(\mathbf{s}^{*})\right)$ $\leq\frac{1}{T}\int_{0}^{T}\left(\int_{\mathbf{x}\sim q_{1-t}(\mathbf{x})}\|\mathbf{s}(1-t,\mathbf{x})-\nabla\log q_{1-t}(\mathbf{x})\|^{2}\mathrm{d}\mathbf{x}\right)\mathrm{d}t$
$\lesssim\frac{1}{T}\int_{0}^{T}\frac{{\mathcal{R}}^{-2\beta}\log\mathcal{R}}{\sigma^{2}(1-t)}\mathrm{d}t\lesssim{\mathcal{R}}^{-2\beta}\log^{2}\mathcal{R}\lesssim n^{-\frac{2\beta}{d^{*}+2\beta}}\log^{2}n.$
Substituting $\mathcal{R}=\lfloor n^{\frac{1}{d^{*}+2\beta}}\rfloor+1\lesssim n^{\frac{1}{d^{*}+2\beta}}$, $m=n^{\frac{d^{*}+8\beta}{d^{*}+2\beta}}$ and $C_{T}=2\beta$ into (21) and (22), we have
@@
which implies that the statistical error can be bounded as
$\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\mathcal{L}(\widehat{\mathbf{s}})-2\overline{\mathcal{L}}_{\widehat{\mathcal{X}}}(\widehat{\mathbf{s}})+\mathcal{L}(\mathbf{s}^{*})\right)+2\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\overline{\mathcal{L}}_{\widehat{\mathcal{X}}}(\widehat{\mathbf{s}})-\widehat{\mathcal{L}}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}(\widehat{\mathbf{s}})\right)\lesssim n^{-\frac{2\beta}{d^{*}+2\beta}}\log^{19}n.$
Thus, we finally obtain
$\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{x}_{t}}\|\widehat{\mathbf{s}}(1-t,\mathbf{x}_{t})-\nabla\log q_{1-t}(\mathbf{x}_{t})\|^{2}\mathrm{d}t\right)\lesssim n^{-\frac{2\beta}{d^{*}+2\beta}}\log^{19}n.$
The proof is complete. ∎
Appendix D Bound $\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}\left[W_{2}(\widetilde{\pi}_{T}^{L},\widehat{p}_{data}^{*})\right]$
In this section, we bound $\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[W_{2}(\widetilde{\pi}_{T}^{L},\widehat{p}_{data}^{*})]$. We can decompose $\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}\left[W_{2}(\widetilde{\pi}_{T}^{L},\widehat{p}_{data}^{*})\right]$ into two terms:
$\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[W_{2}({\widetilde{\pi}_{T}^{L}},\widehat{p}^{*}_{data})]\leq\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[W_{2}(\widetilde{\pi}_{T}^{L},\pi_{T}^{L})]+\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T}^{L},\widehat{p}_{data}^{*})]$
In the following two subsections, we bound $\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[W_{2}(\widetilde{\pi}_{T}^{L},\pi_{T}^{L})]$ and $\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T}^{L},\widehat{p}_{data}^{*})]$ separately.
D.1 Bound $\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T}^{L},\widehat{p}^{*}_{data})]$
In this subsection, we bound the term $\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T}^{L},\widehat{p}^{*}_{data})]$ and prove Lemma 4.7.
Proof of Lemma 4.7.
$\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T}^{L},\widehat{p}^{*}_{data})]$ can be decomposed into following two terms:
$\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T}^{L},\widehat{p}^{*}_{data})]\leq\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T}^{L},\pi_{T})]+\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T},\widehat{p}^{*}_{data})].$
The first term $\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T}^{L},\pi_{T})]$ satisfies that
$\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T}^{L},\pi_{T})]$ $\leq\mathbb{E}_{\mathcal{Y}}\left(\mathbb{E}\|\mathbf{x}_{T}-\mathbf{x}_{T}\mathbf{I}_{\{\|\mathbf{x}_{T}\|_{\infty}\leq L\}}\|^{2}\right)^{\frac{1}{2}}$
$\leq\mathbb{E}_{\mathcal{Y}}\left(\mathbb{E}\|\mathbf{x}_{T}\|^{2}\mathbf{I}_{\{\|\mathbf{x}_{T}\|_{\infty}>L\}}\right)^{\frac{1}{2}}$
$\leq\mathbb{E}_{\mathcal{Y}}\left(\mathbb{E}\|\mathbf{x}_{T}\|^{4}\cdot\mathbb{P}(\|\mathbf{x}_{T}\|_{\infty}>L)\right)^{\frac{1}{4}}.$
Now, we consider the process
where $\bar{\mathbf{w}}_{t}$ is a standard Brownian motion. Then $\bar{\mathbf{x}}_{t}\sim q(\sqrt{t}\sigma,\mathbf{x})$ and $\bar{\mathbf{x}}_{1-t}$ satisfies the reverse SDE
$\mathrm{d}\mathbf{x}_{t}=\sigma^{2}\nabla\log{q}_{1-t}(\mathbf{x}_{t})\mathrm{d}t+\sigma\mathrm{d}\mathbf{w}_{t}.$
[…]
$\mathbb{E}\|\mathbf{x}_{T}\|^{4}=\mathbb{E}\|\bar{\mathbf{x}}_{0}+\sigma\bar{\mathbf{w}}_{1-T}\|^{4}\lesssim 1+\sigma^{4}(1-T)^{2}d^{*}(d^{*}+2)=\mathcal{O}(1)$
and
$\mathbb{P}\left(\|\mathbf{x}_{T}\|_{\infty}>L\right)$ $=\mathbb{P}\left(\|\bar{\mathbf{x}}_{0}+\sigma\bar{\mathbf{w}}_{1-T}\|_{\infty}>L\right)$
$\leq 2d^{*}\exp\left(-\frac{(L-1)^{2}}{2\sigma^{2}(1-T)}\right)$
Using $e^{-x}\leq\frac{1}{x}$ for $x>0$, we have
$\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T}^{L},\pi_{T})]\lesssim\left[\exp\left(-\frac{(L-1)^{2}}{4\sigma^{2}(1-T)}\right)\right]^{\frac{1}{2}}\lesssim\sqrt{1-T}\lesssim n^{-\frac{\beta}{d^{*}+2\beta}}.$
The second term satisfies that
$\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T},\widehat{p}_{data}^{*})]\leq\mathbb{E}_{\mathcal{Y}}\left(\mathbb{E}\|\mathbf{x}_{T}-\mathbf{x}_{1}\|^{2}\right)^{\frac{1}{2}}\leq\mathbb{E}_{\mathcal{Y}}\left(\|\sigma\bar{\mathbf{w}}_{1-T}\|^{2}\right)^{\frac{1}{2}}\lesssim\sqrt{1-T}\lesssim n^{-\frac{\beta}{d^{*}+2\beta}}.$
Therefore, combining the above two inequalities, we finally obtain
$\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T}^{L},\widehat{p}^{*}_{data})]\lesssim n^{-\frac{\beta}{d^{*}+2\beta}}.$ (23)
The proof is complete. ∎
D.2 Bound $\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[W_{2}(\widetilde{\pi}_{T}^{L},\pi_{T}^{L})]$
In this subsection, we bound the term $\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[W_{2}(\widetilde{\pi}_{T}^{L},\pi_{T}^{L})]$. Since $\widetilde{\mathbf{x}}_{T}^{L}$ and $\mathbf{x}_{T}^{L}$ are bounded by $L$, $\widetilde{\pi}_{T}^{L}$ and $\pi_{T}^{L}$ are supported on bounded region $[-L,L]^{d^{*}}$. Then, we obtain the inequality $\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[W_{2}(\widetilde{\pi}_{T}^{L},\pi_{T}^{L})]\lesssim\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[\mathrm{TV}(\widetilde{\pi}_{T}^{L},\pi_{T}^{L})]\lesssim\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[\mathrm{TV}(\widetilde{\pi}_{T},\pi_{T})]$. Therefore, we only need to bound $\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[\mathrm{TV}(\widetilde{\pi}_{T},\pi_{T})]$. We first introduce the following lemma.
Lemma D.1 (Proposition D.1 in [OAS23]).
Let $p_{0}$ be any probability distribution, and $\mathbf{z}=(\mathbf{z}_{t})_{t\in[0,T]}$, $\mathbf{z}^{\prime}=(\mathbf{z}^{\prime}_{t})_{t\in[0,T]}$ be two different processes satisfying
$\mathrm{d}\mathbf{z}_{t}=\mathbf{b}(t,\mathbf{z}_{t})\mathrm{d}t+\sigma(t)\mathrm{d}\mathbf{w}_{t},\leavevmode\nobreak\ \mathbf{z}_{0}\sim p_{0},$
$\mathrm{d}\mathbf{z}^{\prime}_{t}=\mathbf{b}^{\prime}(t,\mathbf{z}^{\prime}_{t})\mathrm{d}t+\sigma(t)\mathrm{d}\mathbf{w}_{t},\leavevmode\nobreak\ \mathbf{z}^{\prime}_{0}\sim p_{0}.$
We define the distributions of $\mathbf{z}_{t}$ and $\mathbf{z}^{\prime}_{t}$ as $p_{t}$ and $p^{\prime}_{t}$, and the path measures of $\mathbf{z}$ and $\mathbf{z}^{\prime}$ as $\mathbb{P}$ and $\mathbb{P}^{\prime}$, respectively. Suppose that the Novikov’s condition holds, i.e.,
$\mathbb{E}_{\mathbb{P}}\left[\exp\left(\frac{1}{2}\int_{0}^{T}\frac{\|\mathbf{b}(t,\mathbf{z}_{t})-\mathbf{b}^{\prime}(t,\mathbf{z}_{t})\|^{2}}{\sigma^{2}(t)}\mathrm{d}t\right)\right]<+\infty.$ (24)
Then, the Radon-Nikodym derivative of $\mathbb{P}^{\prime}$ with respect to $\mathbb{P}$ is
$\frac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{P}}=\exp\left(-\int_{0}^{T}\frac{\mathbf{b}(t,\mathbf{z}_{t})-\mathbf{b}^{\prime}(t,\mathbf{z}_{t})}{\sigma(t)}\mathrm{d}\mathbf{w}_{t}-\int_{0}^{T}\frac{\|\mathbf{b}(t,\mathbf{z}_{t})-\mathbf{b}^{\prime}(t,\mathbf{z}_{t})\|^{2}}{2\sigma^{2}(t)}\mathrm{d}t\right),$
and therefore we have
$\mathrm{KL}(p_{T}|p_{T}^{\prime})\leq\mathrm{KL}(\mathbb{P}|\mathbb{P}^{\prime})=\mathbb{E}_{\mathbb{P}}\left[\frac{1}{2}\int_{0}^{T}\frac{\|\mathbf{b}(t,\mathbf{z}_{t})-\mathbf{b}^{\prime}(t,\mathbf{z}_{t})\|^{2}}{\sigma^{2}(t)}\mathrm{d}t\right].$
Based on Lemma D.1, we can prove Theorem 4.6.